OpenStudy (anonymous):

In a roller-skating rink, the skaters all move at different speeds. Most skate slightly faster or slower than others, but sometimes an adventuresome skater accelerates to an "outlying" speed that is much higher than that of everyone else, creating a hazard. As a result, skating rinks often post a sign that says: Any skater moving faster than the average speed of all the skaters will be asked to leave the rink. It is clear what the sign means, but it is statistically nonsensical. If there are 200 skaters in a rink, all skating at slightly varying speeds, what is the best estimate for how m

OpenStudy (anonymous):

A. 25 B. 0 C. 50 D. 100 E. 75

OpenStudy (anonymous):

oops, the last paragraph is incomplete. It is clear what the sign means, but it is statistically nonsensical. If there are 200 skaters in a rink, all skating at slightly varying speeds, what is the best estimate for how many of them are moving faster than the average speed of all the skaters?

OpenStudy (anonymous):

there

OpenStudy (anonymous):

methinks it might b 25. but thats just a guess

OpenStudy (amistre64):

im guessing d. 100

OpenStudy (amistre64):

the midrange of 200 is 100 ...

OpenStudy (anonymous):

amistre how did u get that?

OpenStudy (anonymous):

amistre ur a freaking genius :P

OpenStudy (amistre64):

im assuming the speeds are normally distributed and that .5 is the probablitity of less then the mean and .5 the probability of above the mean; 200*.5 = 100 ... just a guess tho

OpenStudy (amistre64):

suppose we test this on 5 skaters; 2,3,6,9,13 are the speeds; if we select all the samples of 2, with replacement we get: \begin{array}c 2,2&3,2 &6,2 &9,2 &13,2 &\\ 2,3&3,3 &6,3 &9,3 &13,3 &\\ 2,6&3,6 &6,6 &9,6 &13,6 &\\ 2,9&3,9 &6,9 &9,9 &13,9 &\\ 2,13&3,13 &6,13 &9,13 &13,13 &\\ \end{array} the means of these are: 2.......1 ; 2 2.5 ...2 ; 5 3 ..... 1 ; 3 4 ......2 ; 8 4.5....2 ; 9 5.5....2 ;11 6 ..... 3 ;18 7.5....4 ;30 8 ......2 ;16 9 ..... 1 ; 9 9.5....2 ;19 11 ....2 ;22 13.....1 ;13 ----- 165/25 = 6.6 for a mean 2+3+6+9+13 ------------- = 6.6 for a mean :) 5 6.6 is our mean; and we have 2,3,6 below; and 9,13 above ...so its about half and half was the sample size gets larger